A bouncy, elastic ball of mass 1.20 kg moving with v = 5.70 m/s hits the wall at an angle of 30° and bounces off. If the collision lasted for 0.01 s, what is the average force exerted by the wall on the ball?
Answer #1

we have a relation . imputse I = FAL i = 90-30 = 68 At i moment of ball p= mv = 1.2069 x 5.70 m/s = 6.84 kgols component of m